OpenStudy (anonymous):

Line Q is represented by the following equation: 2x + y = 11 Which equation completes the system that is satisfied by the solution (3, 5)? x + 2y = 15 x + y = 15 x + y = 8 x − y = 2

OpenStudy (lynfran):

is all the info

OpenStudy (anonymous):

yas

OpenStudy (mathstudent55):

Try the point in the given equation and each of the choices. The equation of the choices that has that point as a solution is it.

OpenStudy (mathstudent55):

Whichever equation of the choices has the given point as a solution is the answer.

OpenStudy (anonymous):

can someone just tell me the answer pls

OpenStudy (lynfran):

yes so lets see x+2y=15 3+2(5)=13 so not a solution

OpenStudy (lynfran):

x+y=15 3+5=8 so not a solution

OpenStudy (lynfran):

x+y=8 3+5=8 there u go...lol